NOAA Office of Satellite and Product Operations

NOAA 14 AVHRR Subsystem Summary

Component Status:

Component
Description
Status
BASEPLATEAVHRR baseplateGreen
CHANNEL 1Spectral Channel #1, 0.58 - 0.68 micro metersGreen
CHANNEL 2Spectral Channel #2, 0.725 - 1.0 micro metersGreen
CHANNEL 3Spectral Channel #3, 10.3 to 11.3 micro metersGreen
CHANNEL 4Spectral Channel #4, 3.55 to 3.93 micro metersGreen
CHANNEL 5Spectral Channel #5, 11.5 to 12.5 micro metersGreen
ELECTRONICSAVHRR ElectronicsGreen
IMUInertial Measurement Unit 
OPTICSAVHRR Optical mirrors and telescopesGreen
PSUPower Supply UnitGreen
PTC 1APitch Torque Coil 1A 
PTC 1BPitch Torque Coil 1B 
PTC 2BPitch Torque Coil 2B 
RADIANT COOLERAVHRR Radiant CoolerGreen
SCAN MOTORAVHRR Scan MotorYellow

Subsystem History:

Date Time Component Description
09/28/200617:01:00 Major spacecraft attitude perturbation resulted in autonomous battery undervoltage safing macro implementation turning off AVHRR (scan motor remained on).
08/14/200615:25:00CHANNEL 3Turned AVHRR channels 3,4,5 back ON.
08/10/200617:55:00CHANNEL 5AVHRR IR channels 3, 4 and 5 all commanded off for safing reasons due to high patch temperatures observed during the pitch over maneuver.
08/10/200617:55:00CHANNEL 4AVHRR IR channels 3, 4 and 5 all commanded off for safing reasons due to high patch temperatures observed during the pitch over maneuver.
08/10/200617:55:00CHANNEL 3AVHRR IR channels 3, 4 and 5 all commanded off for safing reasons due to high patch temperatures observed during the pitch over maneuver.
12/20/200516:00:00SCAN MOTORScan motor nominal for 10 days straight. Imagery clean. Scan motor component and AVHRR subsystem both upgraded to YELLOW.
12/09/200500:00:00SCAN MOTORAVHRR scan motor returned to nominal operating range with no bar code imagery. MIRP continues to be rephased once per orbit over the South Pole.
09/29/200516:00:00SCAN MOTORN14 AVHRR downgraded to RED status because of elevated scan motor current and persistent bar coded imagery.
09/27/200501:00:00SCAN MOTORN14 AVHRR scan motor current surged to 280 mAmps (from 150 mA) which WAS accompanied by a concurrent scan motor temp surge. Scan motor current remained elevated this time but appeared to be slowly trending lower (into a more nominal values). Imagery bar codes persist.
09/24/200511:20:07SCAN MOTORN14 AVHRR scan motor current surged to 280 mAmps (from 140 mA) before returning to more nominal values 32 minutes later. Current surge was NOT acompanied by scan motor temp surge. Imagery bar codes started at that time and have continued to be observed as of 26 Sept at 02:00z.
07/21/200512:00:00SCAN MOTORN14 AVHRR scan motor and imagery nominal for the past 5 days. Status upgraded to YELLOW
07/16/200514:30:00SCAN MOTORN14 AVHRR returned to nominal scan motor current and temp. Synch delta also returned to nominal.
04/14/200418:36:00SCAN MOTORToggle AVHRR scan motor OFF then back on 30 seconds later for testing purposes. After turning back on motor current went to pegged levels.
05/01/200300:00:00SCAN MOTORAVHRR scan motor current is consistently high (for over 1 week) and still increasing while image data is consistently degraded. Status downgraded to RED.
03/25/200312:00:00SCAN MOTORAVHRR scan motor no longer showing a problem - status upgraded to yellow.
01/06/200300:00:00SCAN MOTORAVHRR status downgraded back to Red, data quality poor, motor current high
12/26/200200:00:00SCAN MOTORAVHRR Scan Motor Current improved, image quality does not reflect improvement
07/25/200201:00:00SCAN MOTORAVHRR scan motor deteriorating further, data quality poor
06/23/200221:05:00SCAN MOTORAVHRR Scan Motor Current runup.
05/30/200222:00:00SCAN MOTORAVHRR scan motor current surge
01/18/200201:00:00SCAN MOTORAVHRR self corrected. Image quality nominal, currents nearly nominal
11/24/200116:24:00SCAN MOTORAVHRR Scan motor maximum current of 306 ma. Black Body 3,4,5 temperatures became anomalous.
10/18/200119:55:00SCAN MOTORAVHRR high motor currents and synch deltas
10/07/199700:00:00SCAN MOTORSmall increases in the AVHRR scan motor current and synch-delta were observed.
12/30/199410:36:00CHANNEL 5On Orbit
12/30/199410:36:00SCAN MOTOROn Orbit
12/30/199410:36:00ELECTRONICSOn Orbit
12/30/199410:36:00RADIANT COOLEROn Orbit
12/30/199410:36:00OPTICSOn Orbit
12/30/199410:36:00BASEPLATEOn Orbit
12/30/199410:36:00CHANNEL 1On Orbit
12/30/199410:36:00CHANNEL 2On Orbit
12/30/199410:36:00CHANNEL 4On Orbit
12/30/199410:36:00PSUOn Orbit
12/30/199410:36:00CHANNEL 3On Orbit
00:00:00 Toggle MIRP/AVHRR Rephase Enabled the Disabled
00:00:00 Toggle MIRP/AVHRR Rephase Enabled then Disabled.